5x-3y=30
find x and y intercepts

Answer by checkley75(3666)

SETTING X=0 WE GET: 5*0-3Y=30 OR -3Y=30 OR Y=30/-3 OR Y=-10 ANSWER. (0,-10)
SETTING Y=0 WE GET: 5X-3*0=30 OR 5X=30 OR X=30/5 OR X=6 ANSWER. (6,0)